OpenStudy (anonymous):

find the exact function value of cos (-135°)

OpenStudy (anonymous):

cos(-135)=cos(135) since cosX is an even function. but cos(135)=-cos(45) as it is in quadrant 2 of the unit circle

OpenStudy (anonymous):

but we know cos45 as it is a well known working angle. cos45=sqrt(2)/2 therefore -cos45 is -sqrt(2)/2